OpenStudy (anonymous):

Find the fourth roots of 256(cos 240° + i sin 240°).

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

@whpalmer4 Please anybody

OpenStudy (anonymous):

Had this same question up for 30 min and nobody replied :(

OpenStudy (whpalmer4):

do you know De Moivre's Theorem?

OpenStudy (anonymous):

No but i can pull it up real quick if i need it

OpenStudy (anonymous):

[r(costheta + isintheta)] right?

OpenStudy (loser66):

= r^n(cos ntheta + isin n theta) that the formula.

OpenStudy (loser66):

take a look at Demoivre theorem at http://www.intmath.com/complex-numbers/7-powers-roots-demoivre.php

OpenStudy (whpalmer4):

Are you supposed to take the 4th power, or the 4th root?

OpenStudy (anonymous):

OK. Am i supposed to simplify the equation that u gave me? And the only information they give me is the fourth roots of it. Idk about if it wants powers

OpenStudy (anonymous):

take the fourth root of 256 and divide the angle by 4 that is all

OpenStudy (loser66):

do as satellite73 says, kjuchiha.

OpenStudy (anonymous):

fourth root of 256 = 4, divided by 4 = 1 That cant be it, can it?

OpenStudy (anonymous):

the fourth root of 256 is 4 \(240\div 4=60\)

OpenStudy (anonymous):

no, don't divide the fourth root of 256 by 4, divide the angle by 4

OpenStudy (anonymous):

4(cos60 + i*sin60)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wasn't that easy?

OpenStudy (anonymous):

Holy... thank you! So much easier than I thought! took an hour of me not being able to figure it out

OpenStudy (anonymous):

Thanks satellite and all others!

OpenStudy (whpalmer4):

well, interesting that Mathematica gets a different answer...

OpenStudy (anonymous):

Hmm... is that bad?

OpenStudy (anonymous):

now if you want to be fancy, you can say \(\cos(60)=\frac{1}{2}\) and \(\sin(60)=\frac{\sqrt{3}}{2}\) and so the FINAL ANSWER would be \[4\left(\frac{1}{2}+i\frac{\sqrt{3}}{2}\right)\]\[=2+2\sqrt{3}i\]

OpenStudy (anonymous):

don't forget that there are not one, but 4 fourth roots

OpenStudy (anonymous):

What does that mean satellite? there's more?

OpenStudy (anonymous):

we found one of them, lets find another

OpenStudy (loser66):

@satellite73 so, if I apply Demoivre theorem with power of 1/4 , I get the same answer, right?

OpenStudy (anonymous):

the number out front will stay at 4, but we can go around the circle one more time and say \(240+360=600\) and \(600\div 4=150\)

OpenStudy (anonymous):

@Loser66 yes, this is de moivre

OpenStudy (loser66):

bingo.

OpenStudy (anonymous):

360/4 = 90, so shouldnt I just add 90 to each angle? 4(cos 150 + i sin 150) 4(cos 240 + i sin 240) 4(cos 330 + i sin 330)

OpenStudy (anonymous):

so another answer is \[4\left(\cos(150)+i\sin(150)\right)\]

OpenStudy (anonymous):

So all in all its 4(cos 60 + i sin 60) 4(cos 150 + i sin 150) 4(cos 240 + i sin 240) 4(cos 330 + i sin 330)

OpenStudy (anonymous):

yes, add 90 works as well

OpenStudy (anonymous):

you found the sound

OpenStudy (anonymous):

Wow, thats too easy!

OpenStudy (anonymous):

yes, that is kind of the point of demoivre it is easy not hard otherwise, why would you write a number in the form \(a+bi\) as \[r\left(\cos(\theta)+i\sin(\theta)\right)\]?

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw

OpenStudy (whpalmer4):

Good point about the additional roots — I was thrown off by the after that I only got one when I evaluated it, and normally I get multiple roots from Mathematica. Difference was that I hadn't asked it to solve this time, so it didn't present all the candidates.

OpenStudy (anonymous):

i wonder what happens if we do this

OpenStudy (whpalmer4):

\[\{\{x\to -3.4641+2. i\},\{x\to -2.-3.4641 i\},\]\[\{x\to 2.\, +3.4641 i\},\{x\to 3.4641\, -2. i\}\}\]

OpenStudy (whpalmer4):

WA does the same if we ask it to solve x^4 = 256cis(240 deg) http://www.wolframalpha.com/input/?i=%28256%28cos%28240%29%2Bisin%28240%29%29%29%3Dx%5E4